A retired couple has $170,000 to invest to obtain annual income. They want some of it invested in safe Certificates of Deposit yielding 7%. The rest they want to invest in AA bonds yielding 12% per year. How much should they invest in each to realize exactly $17,900 per year?
Answer by mananth(16949)

CD 7.00% per annum ------------- Amount invested =x
AA bonds 12.00% per annum ------------ Amount invested = y
170000
Interest----- 17900

CD 7.00% per annum ---x
AA bonds 12.00% per annum ---y
Total investment
x + 1 y= 170000 -------------1
Interest on both investments
7.00% x + 12.00% y= 17900
Multiply by 100
7 x + 12 y= 1790000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -1190000.00
Add this to (2)
0 x 5 y= 600000
divide by 5
y = 120000
CD 7.00% $ 50000
AA bonds 12.00% $ 120000

CHECK
50000 --------- 7.00% ------- 3500.00
120000 ------------- 12.00% ------- 14400.00
Total -------------------- 17900.00

m.ananth@hotmail.ca